### v3

# On the Unknowable Limits to Prediction

On the Unknowable Limits to Prediction
Jiani Yan For Correspondence: Jiani Yan and Charles Rahal, Leverhulme Centre for Demographic Science, Demographic Science Unit, University of Oxford, OX1 1JD, United Kingdom. Tel: 01865 286170. Email: jiani.yan@sociology.ox.ac.uk and charles.rahal@demography.ox.ac.uk. Acknowledgements: Funding is gratefully acknowledged from the Leverhulme Trust (Grant RC-2018-003) for the Leverhulme Centre for Demographic Science, Nuffield College, and the Economic and Social Research Council. Comments gratefully received from Kyla Chasalow, Mark Verhagen, Ridhi Kashyap, Ben Domingue, Michael Biggs, Sander Wagner, Haohao Lei and Edith Darin. We are especially grateful to comments received from Nicholas Irons. Leverhulme Centre for Demographic Science, University of Oxford Department of Sociology, University of Oxford Wolfson College, University of Oxford Both authors contributed jointly to this work. Charles Rahal Leverhulme Centre for Demographic Science, University of Oxford Nuffield College, University of Oxford Both authors contributed jointly to this work.
(December 7, 2024)
Keywords: Computational Science, Machine Learning, Philosophy of Prediction
We may regard the present state of the universe as the effect of its past and the cause of its future. An intellect which at a certain moment would know all forces that set nature in motion, and all positions of all items of which nature is composed, if this intellect were also vast enough to submit these data to analysis, …nothing would be uncertain.
laplace1814, Essai philosophique sur les probabilités
The classic dichotomisation of prediction error into statically defined ‘reducible’ and ‘irreducible’ terms is unable to capture the intricacies of progressive research design where specific types of truly reducible error are being rapidly eliminated at differential speeds. While this is generalisably true – including in the domain of Large Language Models (LLMs) – it is none more apparent than in social systems where questions regarding the ‘predictability’ of outcomes are gradually reaching the fore (rahal2024rise). Canonical work which convenes common predictive tasks involving conventional social surveys (salganik2020measuring) shows low power, but rapidly emerging computational approaches which utilize non-standard administrative data and information contained within things such as children’s autobiographical essays – often combined with generative Artificial Intelligence and distributed computing (wolfram2022short; savcisens2024using) – begin to allow high accuracy. We propose an enhanced framework which builds upon existing work (hullermeier2021aleatoric) in a way which we believe helpfully decomposes various types of truly reducible ‘epistemic’ error which are in theory eliminable (resulting from a lack of knowledge), and isolates residual, ‘aleatoric’ error (inherent randomness which can never be modeled) that research endeavor will not be able to capture in the limit of human progress. We further emphasize the impossibility of claiming whether things are ‘predictable’, not least in the context of open, dynamically evolving social systems.
Decomposing prediction error into reducible and irreducible terms is not new. [^1] See, for example, Sections 2.9 and 7.3 of hastie2009elements; a landmark instruction manual in the field. This foundational text uses the term ‘reducible’ to taxonomically classify errors into a type that can be eliminated, and the term ‘irreducible’ to denote errors which cannot be eliminated given current information on target variables and feature sets. However, statements regarding ‘predictability’ and ‘irreducibility’ require the reinforcement of an important quantifier; they are entirely conditional on information sets. For example, the insightful mixed-methods work of lundberg2024origins begins its abstract with ‘Why are some life outcomes difficult to predict?’. In that case, they are difficult to predict based on information which is currently measurable, and the ‘task’ at hand. An outcome which is difficult to predict with one feature set may not be ‘unpredictable’ in an aleatoric sense: life trajectories may be trivial to predict in the future. More reflexive terminology is essential, lest conclusions be drawn that there is no space for future positive policy interventions based on ever increasingly accurate predictions which prevent negative outcomes. It is impossible to know whether we have eliminated all reducible epistemic error to approach the practical ‘ceiling’ of accuracy, make statements about how predictable outcomes will become, or understand where we are in the process. Predictability of an outcome can only be asserted when all epistemic errors are eliminated: when ‘unmeasured’ information is measured and constructed perfectly, features are constructed perfectly, and when functional forms are able to be exactly recovered.
Consider Equation 1 which better conceptually decomposes error when predicting within contexts under the assumption of stationarity, where $y_{\text{true}}$ represents the true outcome, measured after constructing on the phenomenon of interest:
$y_{\text{true}}$ $=\underbrace{f^{*}(\mathbf{x}_{\text{true}})}_{\begin{subarray}{c}\text{Predictive}\\ \text{Ceiling}\end{subarray}}+\underbrace{\varepsilon}_{\begin{subarray}{c}\text{Aleatoric}\\ \text{Error}\end{subarray}}$ (1)
$=\underbrace{[f^{*}(\mathbf{x}_{\text{true}})-f(\mathbf{x}_{\text{true}}\mid y_{\text{true}})]}_{\begin{subarray}{c}\text{Model approximation gain}\\ \text{(Epistemic)}\end{subarray}}$
$\quad+\underbrace{[f(\mathbf{x}_{\text{true}}\mid y_{\text{true}})-f(\mathbf{x}_{\text{true}}\mid y_{\text{observed}})]}_{\begin{subarray}{c}\text{Measurement gain from }y\\ \text{(Epistemic)}\end{subarray}}$
$\quad+\underbrace{[f(\mathbf{x}_{\text{true}}\mid y_{\text{observed}})-f(\mathbf{x}_{\text{observed}}\mid y_{\text{observed}})]}_{\begin{subarray}{c}\text{Measurement gain from }\mathbf{x}\\ \text{(Epistemic)}\end{subarray}}$
$\quad+\underbrace{y_{\text{predicted}}}_{\begin{subarray}{c}\text{Current Prediction}\end{subarray}}+\underbrace{\varepsilon}_{\begin{subarray}{c}\text{Irreducible}\\ \text{(Aleatoric)}\end{subarray}}.$
Figure 1: Decomposed Learning Curves: The functional form of learning curves can vary depending on the targets, algorithms, and features being used, and different amounts of truly irreducible error may exist for different target variables (see Supplementary Information Figure S1). As relevant information accumulates, the learning curve is expected to be monotonic, but may not be entirely continuous. Panel ‘a.’ represents a baseline scenario which positions specific studies against how well information on target and features has been measured (and how well such information has been mapped into learning algorithms), while Panels ‘b.’ and ‘c.’ represent changes (and gains and changes in predictive performance) with regards to construct validity of $y$ and $\mathbf{x}$ respectively. Note: ‘obs.’ denotes ‘observed’ variables as per Equation 1.
where $\mathbf{x}_{\text{true}}$ (from a possibly uncountable, infinite collection of features in the universe) and $y_{\text{true}}$ are enhanced information sets over and above $\mathbf{x}_{\text{observed}}$ and $y_{\text{observed}}$ which could tautologically reduce what may currently have been termed ‘irreducible’. [^2] For class labeling tasks (such as binary classification), our conceptual framework remains viable at the aggregate level with appropriately defined model evaluation metrics: we leave this – as well as extensions to things such as probability distributions, entropy, and stochastic processes adapted to filtration – for further work. Predictive accuracy could also improve through better construct validity (o1998empirical); the extent to which the model’s predictions or features accurately capture the theoretical construct or concept the model is intended to measure or represent. [^3] Note: better construction of the target and features may not necessarily involve the collection of additional information, but simply a re-framing of a variable or new utilization of existing information, as depicted in Figures 1b.-c. Incorporation of previously unmeasured information (from $\mathbf{x}_{\text{observed}}$ to $\mathbf{x}_{\text{true}}$) or better measurement of existing features (from $\mathbf{x}_{\text{observed}}$ and $y_{\text{observed}}$ to $\mathbf{x}_{\text{true}}$ and $y_{\text{true}}$) will also reduce epistemic error, as will reductions in model approximation error (from $f\text{ to }f^{*}$). [^4] Kindly see our Supplementary Information for a fully delineated decomposition of Equation 1. To illustrate how these different forms of error reduction work in practice, we visualize our generalized framework in Figure 1. This demonstrates how learning curves evolve as different sources of epistemic uncertainty are reduced (Figure 1a.) and targets and features are constructed differently (Figure 1b.-c.). [^5] Supplementary Information Figure S1a. shows the case where aleatoric error may be low and the speed of improvement relative to information gain is high with large gains from measurement error possible. Supplementary Information Figure S1b. shows the case where aleatoric error may be high, but learning is relatively slower and only minimal gains from reducing learning error are possible.
However, the existence of aleatoric uncertainty means we cannot appreciate a practical limit to prediction, because that itself cannot be measured. We have no reasonable expectation of what data will be collectible, or how powerful deep learning algorithms may become. LLMs – as a specific class of these algorithms – highlight the tension between the scale of available data and the fundamental unpredictability of complex systems. Scaling laws suggest accuracy improves systematically with model size and data, yet within a generalized framework, LLMs probe the boundaries of epistemic and aleatoric uncertainty, revealing where further scaling may yield diminishing marginal returns in the face of inherent unpredictability.
It is only recently that the concept of measuring epigenetic data – let alone interpreting it at scale – has become viable due to advances in molecular biology. Aleatoric error may be large and nontrivial, with suggestions that predictive performance may be bounded well below deterministic precision. Some data may be potentially impossible to collect, and some outcomes may be genuinely stochastic. The ‘time to event’ will also affect how large aleatoric error is; it is easier to predict how somebody will vote at the next election than how the same person will vote in three election’s time. Similarly, re-framing the outcome from ‘when will somebody have a child’ to ‘whether somebody will have a child’ will also affect the size of aleatoric error. We may be able to better approximate ‘perfect prediction’ in some easily constructed and perfectly measurable outcomes (e.g., text prediction and machine translation) than others (e.g., behavioral traits). Research is required to perturb information sets across individual prediction exercises (i.e., comparing better measurement and construct through inclusion/omission), illuminating how quickly each type of epistemic uncertainty is reduced. Work on systematically indexing predictive accuracy is also of utility, as are designs which seek to maximize external validity. Emerging qualitative approaches which explore predictive errors as a function of pivotal moment’s in people’s lives (lundberg2024origins) are also immensely promising, and represent a more nuanced form of information. Collecting more detailed information also necessitates careful discussions around privacy, compliance, and ethical issues, particularly as LLMs often rely on vast datasets that may inadvertently include proprietary or copyrighted material, raising concerns about data provenance, consent, and the potential for intellectual property violations. However, until we have reason to believe that we have eliminated error through better measurement and construction of features and outputs – in addition to the elimination of learning error – we should only tentatively discuss how predictable things are with current data and technology, and not yet uniformly infer a ‘failure to predict’ or an ‘unpredictability of life outcomes’ (garip2020failure, p. 8235).
Supplementary Information
SI.1 Delineating Epistemic and Aleatoric Errors
In predictive modeling, researchers often aim to predict a phenomenon or concept of interest; they aim to construct a concrete and measurable representation of this as $y_{\text{true}}$. However, due to measurement errors that arise from factors such as subjective assessment, accounting inaccuracies, or sampling bias, researchers are often forced to conceptualize this phenomenon which they want to measure as something similar to $y_{\text{observed}}$. Such measurement errors exist in both feature and target variables; it is essential to understand how these errors propagate through the model and affect predictive accuracy. This supplementary section presents a mathematical framework to link $y_{\text{true}}$ with $y_{\text{observed}}$, highlighting epistemic (reducible) and aleatoric (irreducible) errors in prediction, and then equates it to the familiar bias-variance trade-off.
SI.1.1 Problem Setup
Consider the following definition of the ‘true’ target and features, the observable target and features, and the model which binds them as follows.
1. True Target and Features
Allow us to equate the relationship between the ‘true’ target variable and feature set as:
$y_{\text{true}}\coloneqq f^{*}(\mathbf{x}_{\text{true}})+\varepsilon,$ (S2)
where:
  • $y_{\text{true}}$ is the true outcome, measured after constructing on the phenomenon of interest.
  • $\mathbf{x}_{\text{true}}$ represents the best possible feature set both in terms of quantity and in quality; a perfectly chosen, constructed and measured feature set.
  • $f^{*}()$ is the conceptually true underlying function which maps $\mathbf{x}_{\text{true}}$ onto $y_{\text{true}}$.
  • $\varepsilon$ is aleatoric error (inherent randomness) for $y_{\text{true}}$, with $\mathbb{E}[\varepsilon]=0$ and variance $\operatorname{Var}(\varepsilon)=\sigma_{\varepsilon}^{2}$. It is uncorrelated with $f^{*}(\mathbf{x}_{\text{true}})$. A proof of this is available upon request.
2. Observed Target and Features
Next, allow us to introduce measurement error in $y_{\text{true}}$ and $\mathbf{x}_{\text{true}}$ by relating them to what we can observe:
$y_{\text{observed}}$ $\coloneqq y_{\text{true}}+\delta_{y}=f^{*}(\mathbf{x}_{\text{true}})+\varepsilon+\delta_{y},$ (S3)
$\mathbf{x}_{\text{observed}}$ $\coloneqq\mathbf{x}_{\text{true}}+\delta_{\textbf{x}},$ (S4)
where:
  • $\delta_{y}$ is measurement error in $y$, with $\mathbb{E}[\delta_{y}]=\mu_{\delta_{y}}$ and $\operatorname{Var}(\delta_{y})=\sigma_{\delta_{y}}^{2}$.
  • $\delta_{\mathbf{x}}$ is measurement error in $\mathbf{x}$, with $\mathbb{E}[\delta_{\mathbf{x}}]=\mu_{\delta_{\mathbf{x}}}$ and variance $\operatorname{Var}(\delta_{\mathbf{x}})=\sigma_{\delta_{\mathbf{x}}}^{2}$. $\mathbf{x}_{\text{observed}}$ is a vector with $1\times K$ dimensions, with each $x_{k}\in\mathbf{X}$ for $k=\{1,2,...,K\}$ as each one of the predictors. Note: $\mathbf{x}_{\text{observed}}$ can have different dimensions to $\mathbf{x}_{\text{true}}$.
As $y_{\text{observed}}$ is observed based on a constructed concept, $\delta_{y}$ is a function of how well $y_{\text{true}}$ is measured. $\delta_{\mathbf{x}}$ constitutes measurement and construct error both at the single predictor and aggregate levels. At the level of an individual feature ($\textbf{x}_{k}$), it regards how well the feature is constructed and measured, and at the aggregate level, how well the feature set itself is measured (as a function of individual features $\mathbf{x}_{k}$ and the set $\mathbf{X}$ as a whole). Construct validity of $x_{k}\in\mathbf{X}$ can be achieved with the same amount of underlying information (see Figure 1c). For instance, if $x_{k}$ is calculated by averaging the response of multiple items, construct validity can be improved by finding optimal weights by which to combine the individual predictors. $\delta_{\textbf{x}}$ can be reduced through improved construction of each $x_{k}\in\mathbf{X}$; increased sample representativeness, and conceptualizing the feature set so that it contains the most relevant set of predictors. In general, we have assumed that $y_{\text{observed}}$ captures all aleatoric errors in $y_{\text{true}}$; changing $\delta_{y}$ does not impact the level of aleatoric error $\varepsilon$. A discussion of the violation of this condition can be found in SI.1.5
3. Predictive Model
The predictive model is estimated using observable training data $\textbf{x}_{\text{observed}}$ and $y_{\text{observed}}$:
$y_{\text{pred}}\coloneqq f(\textbf{x}_{\text{observed}}\mid y_{\text{observed}}),$ (S5)
where ${f}$ is the estimated model trained on $\{\textbf{x}_{\text{observed}},y_{\text{observed}}\}$. The notation ${f}(\textbf{x}_{\text{observed}}\mid y_{\text{observed}})$ emphasizes that ${f}$ depends on $y_{\text{observed}}$ in training processes. The difference between the conceptually true underlying function $f^{*}()$ and the currently selected predictive function $f()$ on $\textbf{x}_{\text{true}}$ and $y_{\text{true}}$ is defined as follows:
$f(\textbf{x}_{\text{true}}\mid y_{\text{true}})\coloneqq f^{*}(\textbf{x}_{\text{true}})+\delta_{f},$ (S6)
where $\delta_{f}$ is defined based on the optimal $\textbf{x}_{\text{true}}$ and $y_{\text{true}}$, and $\mathbb{E}[\delta_{f}]=\mu_{\delta_{f}},\operatorname{Var}[\delta_{f}]=\sigma_{\delta_{f}}$. The difference between $f(\textbf{x}_{\text{true}}\mid y_{\text{true}})$ and $y_{\text{pred}}$ attributes to improvement in $\delta_{\textbf{x}}$ and $\delta_{y}$; further details can be found in Equation LABEL:eq:y_pred_decomposition_refined.
SI.1.2 Decomposing $y_{\text{true}}$
Following these aforementioned definitions, we can decompose the output of a predictive machine which predicts $y_{\text{true}}$ as perfectly as possible, and in the process link it to existent predictions generated by $f()$ to create $y_{\text{pred}}$ as follows:
$y_{\text{true}}$ $=f^{*}(\textbf{x}_{\text{true}})+\varepsilon$
$=[f^{*}(\textbf{x}_{\text{true}})-f(\textbf{x}_{\text{true}}\mid y_{\text{true}})]$
$\quad+[f(\textbf{x}_{\text{true}}\mid y_{\text{true}})-f(\textbf{x}_{\text{true}}\mid y_{\text{observed}})]$
$\quad+[f(\textbf{x}_{\text{true}}\mid y_{\text{observed}})-f(\textbf{x}_{\text{observed}}\mid y_{\text{observed}})]$
$\quad+f(\textbf{x}_{\text{observed}}\mid y_{\text{observed}})+\varepsilon$
$=\underbrace{-\delta_{f}}_{\shortstack{\scriptsize\text{Model approximation gain} \\ \scriptsize\text{(Epistemic)}}}$
[…]
$\quad+\underbrace{[f(\textbf{x}_{\text{true}}\mid y_{\text{observed}})-f(\textbf{x}_{\text{true}}+\delta_{\textbf{x}}\mid y_{\text{observed}})]}_{\scriptsize\shortstack{\text{Measurement gain from} $\mathbf{x}$ \\ \text{(Epistemic)}}}$
[…]
Therefore, we can achieve improved predictive performance in comparison to existing predictions based on $\textbf{x}_{\text{observed}}$, $y_{\text{observed}}$, and $f()$. This is made possible through opportunities to eliminate measurement and construction error in $\textbf{x}_{\text{observed}}$ (i.e., reducing $\delta_{\textbf{x}}$), measurement in $y_{\text{observed}}$ (i.e., reducing $\delta_{y}$), and better model approximation $f()$ (i.e., reducing $\delta_{f}$). When $\delta_{\textbf{x}}=\delta_{y}=\delta_{f}=0$, we reach a scenario where the difference between $y_{\text{true}}$ and $y_{\text{predicted}}$ is $\varepsilon$: pure aleatoric error. This essentially represents the ‘upper bound’ of predictive accuracy, occasionally referred to as a ‘predictive ceiling’ (garip2020failure).
SI.1.3 Error Decomposition
For any individual observation, the total prediction error (denoted below as $Error$) is defined as the difference between the predicted and true outcome:
$Error$ $\coloneqq y_{\text{pred}}-y_{\text{true}}$
$={f}(\textbf{x}_{\text{observed}}\mid y_{\text{observed}})-y_{\text{true}}$
$=\left[{f}(\textbf{x}_{\text{observed}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}})\right]$
$\quad+\left[{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{true}}\mid y_{\text{true}})\right]$
$\quad+\left[{f}(\textbf{x}_{\text{true}}\mid y_{\text{true}})-f^{*}(\textbf{x}_{\text{true}})\right]+\left[f^{*}(\textbf{x}_{\text{true}})-y_{\text{true}}\right].$
Substituting for Equations S2-S4, we obtain:
$Error$ $=\underbrace{\left[{f}(\textbf{x}_{\text{observed}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{observed}}-\delta_{\textbf{x}}\mid y_{\text{observed}})\right]}_{\shortstack{\scriptsize\text{Error due to measurement error in $\mathbf{x}$} \\ \\ \scriptsize\text{(Epistemic)}}}$
$\quad+\underbrace{\left[{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}}-\delta_{y})\right]}_{\shortstack{\scriptsize\text{Error due to measurement error in $y$ affecting model estimation}\\ \scriptsize\text{(Epistemic)}}}$
[…]
It then logically follows that predicting the target variable as accurately as can be done in practice (see Section SI.1.1) results in the case where all epistemic errors are eliminated, and only aleatoric error remains: $\text{Error}=y_{\text{pred}}-y_{\text{true}}=\varepsilon$.
SI.1.4 Relationship to Existing Frameworks
Our framework mainly decomposes errors in a conceptually simple way (i.e., up until now, we have been dealing with scalar target outcome observations). We next develop the aggregate level decomposition of the expected sum of squared errors and relate our framework to the bias, variance, and error frameworks of canonical textbooks (e.g., james2013introduction; hastie2009elements) as well as papers (pedro2000unified; james2003variance) in the field of statistical learning. Suppose we have a sample with $N$ observations, for each $y_{\text{true},i}\in\textbf{y}_{\text{true}}$ and $y_{\text{pred},i}\in\textbf{y}_{\text{predicted}}$ for $i=\{1,2,...,N\}$. Equation LABEL:eq:error_decomposition holds as follows:
$Error$ $=y_{\text{pred}}-y_{\text{true}}$
$=\underbrace{\left[{f}(\textbf{x}_{\text{observed}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{observed}}-\delta_{\textbf{x}}\mid y_{\text{observed}})\right]}_{\shortstack{\scriptsize\text{Error due to measurement error in $\mathbf{x}$} \\ \\ \scriptsize\text{(Epistemic)}}}$
$\quad+\underbrace{\left[{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}}-\delta_{y})\right]}_{\shortstack{\scriptsize\text{Error due to measurement error in $y$ affecting model estimation}\\ \scriptsize\text{(Epistemic)}}}$
[…]
$=\delta^{\text{error}}_{\mathbf{x}}+\delta^{\text{error}}_{y}+\delta_{f}+\varepsilon,$ (S10)
where:
  • $\delta^{\text{error}}_{\textbf{x}}\coloneqq{f}(\textbf{x}_{\text{observed}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{observed}}-\delta_{\textbf{x}}\mid y_{\text{observed}})$; the difference in predicted values owing to $\delta_{\textbf{x}}$. $\mathbb{E}[\delta^{\text{error}}_{\textbf{x}}]=\mu^{\text{error}}_{\textbf{x}}$ and variance $\operatorname{Var}(\delta^{\text{error}}_{\textbf{x}})=\sigma^{2}_{\delta^{\text{error}}_{\textbf{x}}}$. As $\delta^{\text{error}}_{\textbf{x}}$ is expressed as the difference between the fitted values, the calculation of $\operatorname{Var}(\delta^{\text{error}}_{\textbf{x}})$ depends on the specific form of $f$, the distribution of $\textbf{x}_{\text{observed}}\text{ and }\textbf{x}_{\text{observed}}-\delta_{\textbf{x}}$, and their dependencies.
  • $\delta^{\text{error}}_{y}\coloneqq{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}}-\delta_{y})$; the difference in predicted values owing to $\delta_{y}$. $\mathbb{E}[\delta^{\text{error}}_{y}]=\mu^{\text{error}}_{y}$ and variance $\operatorname{Var}(\delta^{\text{error}}_{y})=\sigma^{2}_{\delta^{\text{error}}_{y}}$. Similarly, the calculation of $\operatorname{Var}(\delta^{\text{error}}_{y})$ depends on the specific form of $f$, the distribution of $y_{\text{observed}}\text{ and }y_{\text{observed}}-\delta_{y}$, and their dependencies.
  • $\delta_{f}$ is defined as before in Equation S6.
Therefore,
$\mathbb{E}[Error]$ $=\mathbb{E}[\delta^{\text{error}}_{\textbf{x}}+\delta^{\text{error}}_{y}+\delta_{f}+\varepsilon]$
$=\mu^{\text{error}}_{\textbf{x}}+\mu^{\text{error}}_{y}+\mu_{f}.$ (S11)
Since $\varepsilon$ is uncorrelated with other sources of error:
$\operatorname{Var}(Error)$ $=\operatorname{Var}[(\delta^{\text{error}}_{\textbf{x}}+\delta^{\text{error}}_{y}+\delta_{f}+\varepsilon)]$
$=\operatorname{Var}[(\delta^{\text{error}}_{\textbf{x}}+\delta^{\text{error}}_{y}+\delta_{f})]+\operatorname{Var}(\varepsilon).$ (S12)
With Equations S11-S12 we can write our squared error loss as:
$\mathbb{E}[(y_{\text{true}}-y_{\text{pred}})^{2}]$ $=\mathbb{E}[(y_{\text{true}}-y_{\text{pred}})]^{2}+\operatorname{Var}[(y_{\text{true}}-y_{\text{pred}})]$
$=\mathbb{E}[Error]^{2}+\operatorname{Var}(Error)$
$=(\underbrace{\mu^{\text{error}}_{\textbf{x}}+\mu^{\text{error}}_{y}+\mu_{\delta_{f}}}_{\text{Bias}})^{2}+\underbrace{\operatorname{Var}[(\delta^{\text{error}}_{\textbf{x}}+\delta^{\text{error}}_{y}+\delta_{f})]}_{\text{Variance}}+\underbrace{\operatorname{Var}(\varepsilon)}_{\text{AleatoricError}},$ (S13)
which results in the familiar Irreducible $\text{Irreducible Error}+\text{Bias}^{2}+\text{Variance}$ framework akin to Equation 7.9 of hastie2009elements. For simplicity, we don’t expand $\operatorname{Var}[(\delta^{\text{error}}_{\textbf{x}}+\delta^{\text{error}}_{y}+\delta_{f})]$, of which the full expansion contains the covariance terms between each of the $\delta$ terms.
SI.1.5 Partially captured aleatoric error in observed data
In Equations S3-S4, we assume that $y_{\text{observed}}$ captures all aleatoric error in $y_{\text{true}}$. However, since $y_{\text{observed}}$ is measured after constructing $y_{\text{true}}$, this assumption may not hold in all situations. For example, observed data may be a biased sub-sample of the whole population. With such observed data, we can only optimize epistemic errors ($\delta_{f},\delta_{\textbf{x}}\text{ and }\delta_{y}$) that are inherent in training data. In this case – even when all epistemic error is reduced – $\varepsilon_{y_{\text{observed}}}$ cannot represent the true $\varepsilon_{y}$. Consider a true target variable which is measurable across a population ($y_{\text{true}}$), but $y_{\text{observed}}$ only covers a non-representative 70% of the population. In this case, $\varepsilon_{y_{\text{observed}}}$ only captures the $\varepsilon_{y}$ of the observed sample, and lacks the ability to capture the component which would be inherent in the remaining 30% of the population. The representativeness of the error term $\varepsilon_{y_{\text{observed}}}$ depends on how well $y_{\text{observed}}$ represents true reality ($y_{\text{true}}$). It also affects assumptions we need to make in terms of the bias-variance decomposition as per Equations S12-S13. If $\varepsilon_{y_{\text{observed}}}\neq\varepsilon_{y}$, the covariance between $\varepsilon_{y_{\text{observed}}}$ and $\delta_{f},\delta^{\text{error}}_{y}\text{ and }\delta^{\text{error}}_{\textbf{x}}$ are no longer zero. The $\operatorname{Var}(Error)$ term must also incorporate $\operatorname{Cov}(\varepsilon_{y_{\text{observed}}},\delta_{f}),\operatorname{Cov}(\varepsilon_{y_{\text{observed}}},\delta^{\text{error}}_{y}),\operatorname{Cov}(\varepsilon_{y_{\text{observed}}},\delta^{\text{error}}_{\textbf{x}})$. Therefore, guaranteeing the representativeness of $y_{\text{observed}}$ to $y_{\text{true}}$ (which also affects the representativeness of $\textbf{x}_{\text{observed}}$ to $\textbf{x}_{\text{true}}$ from a measurement perspective) is a necessary step before considering epistemic errors in the form of $\delta_{f},\delta_{y},\text{ and }\delta_{\textbf{x}}$, as it interrogates whether the questions we are answering are really what we meant to ask.
Supplementary Figures
Figure S1: Decomposed Learning Curves: Two complimentary example. Panel ‘a.’ represents a target variable with relatively little aleatoric error, and a rapid learning rate with large gains from eliminating measurement error. Panel ‘b.’ represents a target variable with relatively higher aleatoric error, a slower learning rate, and relatively fewer gains from improving measurement. Note: there is no requirement for such learning curves to be continuous functions. Dashed horizontal red lines denote what is commonly known as the ‘predictive ceiling’ (no prediction error). Golden arrows denote the rate at which reducible error is reducing. Dashed blue lines represents a less accurate learning algorithm, and the solid blue line represents the case where learning error is entirely eliminated.
Supplementary References
  • Garip (2020) Garip, F. (2020). What failure to predict life outcomes can teach us. Proceedings of the National Academy of Sciences 117(15), 8234–8235.
  • Hastie et al. (2009) Hastie, T., R. Tibshirani, J. H. Friedman, and J. H. Friedman (2009). The elements of statistical learning Volume 2. Springer.
  • James (2003) James, G. (2003). Variance and bias for general loss functions. Machine learning 51, 115–135.
  • James et al. (2013) James, G., D. Witten, T. Hastie, R. Tibshirani, et al. (2013). An introduction to statistical learning Volume 112. Springer.
  • Pedro (2000) Pedro, D. (2000). A unified bias-variance decomposition and its applications. In 17th International conference on machine learning, pp.  231–238.